On fibrations with Grassmann manifolds as fibers

Július Korbaš

The aim of this talk is to present some results on Z_{2}-cohomology properties of smooth fiber bundles and of Serre fibrations having complex or real Grassmann manifolds as their fibers.

We mainly concentrate on describing situations when the fiber inclusion (into the total space) induces an epimorphism in Z_{2}-cohomology. This is related to other topological questions.

The talk will be based on the paper On fibrations with Grassmannian fibers, Bull. Belgian Math. Soc. 8 (2001), 119-130; more recent results will also be included.

Department of Algebra,

Faculty of Mathematics and Physics, Comenius University, Mlynská dolina, SK-842 48 Bratislava 4, SLOVAKIA (korbas@fmph.uniba.sk)